# Landau.bib

### v1

@@ 2.4 Entropy dissipation and Fisher information
    We multiply (2.26) by $\phi f(y,w)$ and integrate over $(y,w)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$, noting that this will annihilate the last bracket term in the second line of (2.26), to derive
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi q_{ij}f(y,w){\,\rm d}y{\,\rm d}w$
    $=\Big{(}v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big{)}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi f(y,w){\,\rm d}y{\,\rm d}w$
    $+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{j}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
    $+v_{i}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi^{\prime}w_{j}f(y,w){\,\rm d}y{\,\rm d}w-v_{j}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi^{\prime}w_{i}f(y,w){\,\rm d}y{\,\rm d}w.$
    Next we multiply (2.26) by $\phi w_{i}f(y,w)$ and integrate over $(y,w)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ to derive
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi q_{ij}w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
    $=\Big{(}v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big{)}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
    $+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{i}w_{j}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{i}^{2}f(y,w){\,\rm d}y{\,\rm d}w$
    $+v_{i}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi^{\prime}w_{i}w_{j}\phi f(y,w){\,\rm d}y{\,\rm d}w$
    $-v_{j}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}(\phi+w_{i}^{2}\phi^{\prime})f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{j}f(y,w){\,\rm d}y{\,\rm d}w,$
@@
    where to obtain the last term we have used the following calculations
[…]
    $\qquad=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{i}^{2}\phi(w)\partial_{j}f(y,w){\,\rm d}y{\,\rm d}w-\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{i}w_{j}\phi(w)\partial_{i}f(y,w){\,\rm d}y{\,\rm d}w$
    $\qquad=-\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{i}^{2}w_{j}\phi^{\prime}(w)f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}\partial_{i}(w_{i}\phi(w))f(y,w){\,\rm d}y{\,\rm d}w$
    $\qquad=-\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{i}^{2}w_{j}\phi^{\prime}(w)f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}(\phi(w)+w_{i}^{2}\phi^{\prime}(w))f(y,w){\,\rm d}y{\,\rm d}w$
    $\qquad=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}w_{j}\phi(w)f(y,w){\,\rm d}y{\,\rm d}w.$
    Similarly, testing (2.26) by $\phi w_{j}f(y,w)$ yields
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi q_{ij}w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
    $=\Big{(}v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big{)}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
    $+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{j}^{2}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{i}w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
    $-v_{j}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi^{\prime}w_{i}w_{j}\phi f(y,w){\,\rm d}y{\,\rm d}w$
    $+v_{j}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}(\phi+w_{j}^{2}\phi^{\prime})f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi w_{j}f(y,w){\,\rm d}y{\,\rm d}w.$
@@
    We thus obtain a $3\times 3$-system of linear equations for $\Big{(}v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big{)}$, $\frac{\partial_{i}f}{f}(x,v)$ and $\frac{\partial_{j}f}{f}(x,v)$. Solving this system by Cramer’s formula, we get
    $\frac{\partial_{i}f}{f}(x,v)=\frac{\Delta_{q}(f)}{\Delta_{\phi}(f)}.$
    In the above equation $\Delta_{q}(f)$ is given by
    $\Delta_{q}(f)$ $=\operatorname{Det}\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi f(y,w)\begin{pmatrix}1&w_{i}&q_{ij}+P^{ij}_{1}\\ w_{i}&w_{i}^{2}&w_{i}q_{ij}+P^{ij}_{2}\\ w_{j}&w_{i}w_{j}&w_{j}q_{ij}+P^{ij}_{3}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big{)},$
@@
    And $\Delta_{\varphi}(f)$ is given by
    $\Delta_{\phi}(f)$ $=\operatorname{Det}\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi f(y,w)\begin{pmatrix}1&w_{i}&w_{j}\\ w_{i}&w_{i}^{2}&w_{i}w_{j}\\ w_{j}&w_{i}w_{j}&w_{j}^{2}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big{)}.$
    One can closely follow [Des16, Proposition 4] to show that $f\in L^{1}_{2,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ and $|\mathcal{H}(f)|<+\infty$ ensure that $\Delta_{\phi}\geq C>0$. Then we have
    $\Big{|}\frac{\nabla_{v}f}{f}(x,v)\Big{|}\lesssim\sum_{i,j}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{2}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi f(\sum_{k=1}^{3}|P^{ij}_{k}|+|q_{ij}|\langle w\rangle\big{)}{\,\rm d}y{\,\rm d}w$
    $\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{2}\Big{(}\langle v\rangle\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\langle w\rangle^{2}(\phi+\phi^{\prime}){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi f|q|\langle w\rangle{\,\rm d}y{\,\rm d}w\Big{)}$
    $\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{2}\Big{(}\langle v\rangle\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f|q|\langle w\rangle^{-\frac{\gamma}{2}}{\,\rm d}y{\,\rm d}w\Big{)}.$
    Hence, we have
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\langle v\rangle^{\gamma}f(x,v)\Big{|}\frac{\nabla_{v}f}{f}(x,v)\Big{|}^{2}{\,\rm d}x{\,\rm d}v$
    $\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{6}\Big{(}\|f\|_{L^{1}_{0,2+\gamma}(\mathbb{R}^{3}\times\mathbb{R}^{3})}$
    $+\int_{\mathbb{R}^{6}}\langle v\rangle^{\gamma}f(x,v)\big{|}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(y,w)|q|\langle w\rangle^{-\frac{\gamma}{2}}{\,\rm d}y{\,\rm d}w\big{|}^{2}{\,\rm d}x{\,\rm d}v\Big{)}$
    $\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{6}\big{(}\|f\|_{L^{1}_{0,2+\gamma}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\mathcal{D}(f)C^{-}_{-\gamma,-\gamma}(f)\big{)}$
    whereas in (2.23).
    We combine the bounds on Fisher information and the cross-Fisher information (2.25) to derive (2.17), where $C^{-}_{-\gamma,-\gamma}$ is given in (2.28) below and
    $C(f)$ $\leq\|f\|_{L^{1}_{0,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})}^{6}(\|f\|_{L^{1}_{0,2+\gamma}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+C^{-}_{-\gamma,0}(f)).$
@@
Let $\alpha\in(-3,\infty)$ and $\delta\in\mathbb{R}$. Let $f\in\mathcal{P}\cap\mathcal{S}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. The following estimates hold:
@@
    $\|f*_{v}|v-\cdot|^{\alpha}\langle\cdot\rangle^{\delta}\|_{L^{1}(\mathbb{R}^{3})}\leq C^{+}_{\alpha,\delta}(f)\langle v\rangle^{\alpha}\quad\forall v\in\mathbb{R}^{3},$
@@
    $C^{+}_{\alpha,\delta}(f)\lesssim\|\langle v\rangle^{\alpha+\delta_{+}}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})},\quad\delta_{+}=\max\{\delta,0\}.$ (2.27)
  • If $\alpha\in(-3,0)$, then we have
    $\|f*_{v}|v-\cdot|^{\alpha}\langle\cdot\rangle^{\delta}\|_{L^{1}(\mathbb{R}^{3})}\leq C^{-}_{\alpha,\delta}(f)\langle v\rangle^{\alpha}\quad\forall v\in\mathbb{R}^{3},$
@@
    $C^{-}_{\alpha,\delta}(f)\lesssim\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{p}_{v}L^{1}_{x}}$ (2.28)
    for some $p>\frac{3}{3-|\alpha|}$.
@@
The proof follows [carrillo2024landau] with appropriate modification according to the spatial kernel $x$.
@@
  • The case of $\alpha\in(-3,0)$: We consider the cases of $\{|v|\leq 1\}$ and $\{|v|\geq 1\}$ separately.
    In the case of $|v|\leq 1$, we only need to show the boundedness of $\|f*_{v}|v-\cdot|^{\alpha}\langle\cdot\rangle^{\delta}\|_{L^{1}(\mathbb{R}^{3})}$. Notice that
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$ (2.29)
    $=$ ${}\Big{(}\int_{\mathbb{R}^{3}\times\{|v-v_{*}|\geq 1\}}+\int_{\mathbb{R}^{3}\times\{|v-v_{*}|\leq 1\}}\Big{)}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\int_{\mathbb{R}^{3}\times\{|v-v_{*}|\leq 1\}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*},$
    where we have used the inequality $|v-v_{*}|^{\alpha}\leq 1$ (since $|v-v_{*}|\geq 1$ and $\alpha<0$) to estimate the first term. On the domain $\mathbb{R}^{3}\times\{|v-v_{*}|\leq 1\}$, we have
    $\int_{\mathbb{R}^{3}\times\{|v-v_{*}|\leq 1\}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$ (2.30)
    $=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x_{*},v_{*})\langle v_{*}\rangle^{\delta}\chi_{B_{1}}|v-v_{*}|^{\alpha}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $=\big{(}\int_{\mathbb{R}^{3}}\langle\cdot\rangle^{\delta}f(x_{*},\cdot){\,\rm d}x_{*}\big{)}*_{v_{*}}(\chi_{B_{1}}|\cdot|^{\alpha})$
@@
    where $B_{1}=\{v\mid|v|\leq 1\}$ and $\chi_{B_{1}}$ denotes its characteristic function, and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Notice that $p>\frac{3}{3-|\alpha|}$ ensures that $30\leq|\alpha|p^{\prime}<$ 3, and hence $|v|^{\alpha}\in L^{p^{\prime}}(B_{1})$.

### v2

@@ 2.4 Entropy dissipation and Fisher information
We multiply (2.18) by $\phi\kappa(x-y)f(y,w)$ and integrate over $(y,w)\in\mathbb{R}^{2d}$, noting that this will annihilate the last bracket term in the second line of (2.18), to derive
$\int_{\mathbb{R}^{2d}}\phi\kappa q_{ij}f(y,w){\,\rm d}y{\,\rm d}w$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$+v_{i}\int_{\mathbb{R}^{2d}}\phi^{\prime}\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w-v_{j}\int_{\mathbb{R}^{2d}}\phi^{\prime}\kappa w_{i}f(y,w){\,\rm d}y{\,\rm d}w.$ (2.21)
Next we multiply (2.18) by $\phi\kappa w_{i}f(y,w)$ and integrate over $(y,w)\in\mathbb{R}^{2d}$ to derive
$\int_{\mathbb{R}^{2d}}\phi\kappa q_{ij}w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}w_{j}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}^{2}f(y,w){\,\rm d}y{\,\rm d}w$
$+v_{i}\int_{\mathbb{R}^{2d}}\phi^{\prime}w_{i}w_{j}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$-v_{j}\int_{\mathbb{R}^{2d}}(\phi+w_{i}^{2}\phi^{\prime})\kappa f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w,$ (2.22)
@@
where to obtain the last term, we have used the following calculations
$\int_{\mathbb{R}^{2d}}\Big(w_{i}\frac{\partial_{j}f}{f}(y,w)-w_{j}\frac{\partial_{i}f}{f}(y,w)\Big)w_{i}\phi\kappa f(y,w)\,dydw$
$\qquad=\int_{\mathbb{R}^{2d}}w_{i}^{2}\phi\kappa\partial_{j}f(y,w){\,\rm d}y{\,\rm d}w-\int_{\mathbb{R}^{2d}}w_{i}w_{j}\phi\kappa\partial_{i}f(y,w){\,\rm d}y{\,\rm d}w$
$\qquad=-\int_{\mathbb{R}^{2d}}w_{i}^{2}w_{j}\phi^{\prime}\kappa f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}w_{j}\partial_{i}(w_{i}\phi)\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$\qquad=-\int_{\mathbb{R}^{2d}}w_{i}^{2}w_{j}\phi^{\prime}\kappa f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}w_{j}(\phi+w_{i}^{2}\phi^{\prime}\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$\qquad=\int_{\mathbb{R}^{2d}}w_{j}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w.$
Similarly, testing (2.18) by $\phi\kappa w_{j}f(y,w)$ yields
$\int_{\mathbb{R}^{2d}}\phi\kappa q_{ij}w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
$=\Big(v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)\Big)\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
$+\frac{\partial_{i}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}^{2}f(y,w){\,\rm d}y{\,\rm d}w-\frac{\partial_{j}f}{f}(x,v)\int_{\mathbb{R}^{2d}}\phi\kappa w_{i}w_{j}f(y,w){\,\rm d}y{\,\rm d}w$
$-v_{j}\int_{\mathbb{R}^{2d}}\phi^{\prime}\kappa w_{i}w_{j}\phi\kappa f(y,w){\,\rm d}y{\,\rm d}w$
$+v_{j}\int_{\mathbb{R}^{2d}}(\phi\kappa+w_{j}^{2}\phi^{\prime}\kappa)f(y,w){\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}\phi\kappa w_{j}f(y,w){\,\rm d}y{\,\rm d}w.$ (2.23)
@@
From (2.21), (2.22) and (2.23), we obtain a $3\times 3$-system of linear equations for $v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)$, $\frac{\partial_{i}f}{f}(x,v)$ and $\frac{\partial_{j}f}{f}(x,v)$. Solving this system by Cramer’s formula, we get
$\frac{\partial_{i}f}{f}(x,v)=\frac{\Delta_{i}(f)}{\Delta_{\phi}(f)}\quad\text{and}\quad v_{i}\frac{\partial_{j}f}{f}(x,v)-v_{j}\frac{\partial_{i}f}{f}(x,v)=\frac{\Delta_{ij}(f)}{\Delta_{\phi}(f)}.$
In the above equation: $\Delta_{i}(f)$ and $\Delta_{ij}(f)$ are given by
$\Delta_{i}(f)$ $=\operatorname{Det}\Big(\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w)\begin{pmatrix}1&q_{ij}+P^{ij}_{1}&w_{i}\\ w_{i}&w_{i}q_{ij}+P^{ij}_{2}&w_{i}^{2}\\ w_{j}&w_{j}q_{ij}+P^{ij}_{3}&w_{i}w_{j}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big),$
$\Delta_{ij}(f)$ $=\operatorname{Det}\Big(\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w)\begin{pmatrix}q_{ij}+P^{ij}_{1}&w_{i}&w_{j}\\ w_{i}q_{ij}+P^{ij}_{2}&w_{i}w_{j}&w_{i}^{2}\\ w_{j}q_{ij}+P^{ij}_{3}&w_{j}^{2}&w_{i}w_{j}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big),$
@@
And $\Delta_{\phi}(f)$ is given by
$\Delta_{\phi}(f)$ $=\operatorname{Det}\Big(\int_{\mathbb{R}^{2d}}\phi\kappa f(y,w)\begin{pmatrix}1&w_{i}&w_{j}\\ w_{i}&w_{i}w_{j}&w_{i}^{2}\\ w_{j}&w_{j}^{2}&w_{i}w_{j}\end{pmatrix}{\,\rm d}y{\,\rm d}w\Big).$
We define
$\mathbb{F}(x,w)=\int_{\mathbb{R}^{d}}f(y,w)\kappa(x-y){\,\rm d}y.$
By using the convexity of $x\mapsto x\log x$ and Jensen’s inequality, we have $\mathcal{H}(\mathbb{F})\leq\mathcal{H}(f)$. And we have $\|\mathbb{F}\|_{L^{1}_{2,2}(\mathbb{R}^{2d})}\leq C(\|\kappa\|_{L^{\infty}})\|f\|_{L^{1}_{2,2}(\mathbb{R}^{2d})}$. One can closely follow [Des16, Proposition 4] to show that $\mathbb{F}\in L^{1}_{2,2}(\mathbb{R}^{2d})$ and $|\mathcal{H}(\mathbb{F})|<+\infty$ ensure that $\Delta_{\phi}\geq C>0$. Then we have
$\Big|\frac{\nabla_{v}f}{f}(x,v)\Big|\lesssim\sum_{i,j}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{2}\int_{\mathbb{R}^{2d}}\phi\kappa f(\sum_{k=1}^{3}|P^{ij}_{k}|+|q_{ij}|\langle w\rangle\big){\,\rm d}y{\,\rm d}w$
$\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{2}\Big(\langle v\rangle\int_{\mathbb{R}^{2d}}f\langle w\rangle^{2}(\phi+\phi^{\prime})\kappa{\,\rm d}y{\,\rm d}w+\int_{\mathbb{R}^{2d}}\phi\kappa f|q|\langle w\rangle{\,\rm d}y{\,\rm d}w\Big)$
$\lesssim\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{2}\Big(\|\kappa\|_{L^{\infty}}\langle v\rangle\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}+\int_{\mathbb{R}^{2d}}\kappa f|q|{\,\rm d}y{\,\rm d}w\Big).$
Using (2.20), we deduce that the same bounds also hold for $|\Big[v,\frac{\nabla_{v}f}{f}\Big]_{i,j}|$.
Hence, by using Cauchy–Schwarz inequality, we have
$\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}f(x,v)\Big|\frac{\nabla_{v}f}{f}(x,v)\Big|^{2}{\,\rm d}x{\,\rm d}v$
$\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\Big(\|f\|_{L^{1}_{0,2+\gamma_{+}}(\mathbb{R}^{2d})}$
$+\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}f(x,v)\big|\int_{\mathbb{R}^{2d}}\kappa f(y,w)|q|{\,\rm d}y{\,\rm d}w\big|^{2}{\,\rm d}x{\,\rm d}v\Big)$
$\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\Big(\|f\|_{L^{1}_{0,2+\gamma_{+}}(\mathbb{R}^{2d})}$
$+\int_{\mathbb{R}^{2d}}f(x,v)\big(\int_{\mathbb{R}^{2d}}\kappa f(y,w)|q|^{2}|v-w|^{\gamma}{\,\rm d}y{\,\rm d}w\big)$
$\times\Big(\int_{\mathbb{R}^{2d}}\kappa f(y,w)|v-w|^{-\gamma}\langle v\rangle^{\gamma}\Big){\,\rm d}x{\,\rm d}v\Big)$
$\lesssim$ ${}\|f\|_{L^{1}_{0,2}(\mathbb{R}^{2d})}^{6}\big(\|f\|_{L^{1}_{0,2+\gamma_{+}}(\mathbb{R}^{2d})}+\mathcal{D}(f)C^{\pm}_{-\gamma,0}(f)\big)$
where $C^{\pm}_{-\gamma,0}$ are given in (2.25) and (2.24) below. The same bounds hold for the cross Fisher information.
@@
Let $\alpha\in(-d,\infty)$ and $\delta\in\mathbb{R}$. Let $f\in\mathcal{P}\cap\mathcal{S}(\mathbb{R}^{2d})$. The following estimates hold:
@@
    $\|\langle v\rangle^{\delta}f*_{v}|v|^{\alpha}\|_{L^{1}_{x}}\leq C^{+}_{\alpha,\delta}(f)\langle v\rangle^{\alpha}\quad\forall v\in\mathbb{R}^{d},$
@@
    $C^{+}_{\alpha,\delta}(f)\lesssim\|\langle v\rangle^{\alpha+\delta_{+}}f\|_{L^{1}(\mathbb{R}^{2d})},\quad\delta_{+}=\max\{\delta,0\}.$ (2.24)
  • If $\alpha\in(-d,0)$, then we have
    $\|\langle v\rangle^{\delta}f*_{v}|v|^{\alpha}\|_{L^{1}_{x}}\leq C^{-}_{\alpha,\delta}(f)\langle v\rangle^{\alpha}\quad\forall v\in\mathbb{R}^{d},$
@@
    $C^{-}_{\alpha,\delta}(f)\lesssim\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{1}(\mathbb{R}^{2d})}+\min(\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{p}_{v}L^{1}_{x}},\|\langle v\rangle^{|\alpha|+\delta_{+}}f_{t}\|_{L^{1}_{x}L^{p}_{v}})$ (2.25)
    for some $p>\frac{d}{d-|\alpha|}$.
@@
The proof follows [CDDW24] with appropriate modification according to the spatial kernel $x$.
@@
  • The case of $\alpha\in(-d,0)$: We consider the cases of $\{|v|\leq 1\}$ and $\{|v|\geq 1\}$ separately.
    In the case of $|v|\leq 1$, we only need to show the boundedness of $\|f*_{v}|v-\cdot|^{\alpha}\langle\cdot\rangle^{\delta}\|_{L^{1}(\mathbb{R}^{d})}$. Notice that
    $\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$ (2.26)
    $=$ ${}\Big(\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\geq 1\}}+\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}\Big)f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{2d})}+\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*},$
    where we have used the inequality $|v-v_{*}|^{\alpha}\leq 1$ (since $|v-v_{*}|\geq 1$ and $\alpha<0$) to estimate the first term. On the domain $\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}$, we have
    $\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$ (2.27)
    $=\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})\langle v_{*}\rangle^{\delta}\chi_{B_{1}}|v-v_{*}|^{\alpha}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $=\big(\int_{\mathbb{R}^{d}}\langle\cdot\rangle^{\delta}f(x_{*},\cdot){\,\rm d}x_{*}\big)*_{v_{*}}(\chi_{B_{1}}|\cdot|^{\alpha})$
@@
    where $B_{1}=\{v\mid|v|\leq 1\}$ and $\chi_{B_{1}}$ denotes its characteristic function, and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Notice that $p>\frac{d}{d-|\alpha|}$ ensures that $0\leq|\alpha|p^{\prime}<d$, and hence $|v|^{\alpha}\in L^{p^{\prime}}(B_{1})$. Notice that, in the above, if we first apply Young’s inequality then integrate of $x_{*}$, then we obtain the bound in terms of $\|\langle v\rangle^{\delta}f\|_{L^{1}_{x}L^{p}_{v}}$. This argument also applies to the following estimates.